OpenStudy (anonymous):

Consider the system of three linear equations involving three variables. a_1 x+b_1 y+c_1 z=k_1 a_2 x+b_2 y+c_2 z=k_2 a_3 x+b_3 y+c_3 z=k_3 Where a_1,a_2,a_3,b_1,b_2,b_3,c_1,c_2,c_3,k_1,k_2,and k_3 are real numbers. Write these equation in matrix form. State creamers rule for solving for x, y and z in terms of these real numbers

OpenStudy (unklerhaukus):

\[\begin{align}a_1x+b_1y+c_1z&=k_1\\ a_2x+b_2y+c_2z&=k_2\\ a_3x+b_3y+c_3z&=k_3\end{align}\] \[\begin{bmatrix}a_1&b_1&c_1\\a_2&b_2&c_2\\a_3&b_3&c_3\end{bmatrix}\begin{bmatrix}x\\y\\z\end{bmatrix}=\begin{bmatrix}k_1\\k_2\\k_3\end{bmatrix}\] \[\left[\begin{array}{ccc|c}a_1&b_1&c_1&k_1\\a_2&b_2&c_2&k_2\\a_3&b_3&c_3&k_3\end{array}\right]\]

OpenStudy (anonymous):

is that all?

OpenStudy (unklerhaukus):

i dont remember what creamers rule is

OpenStudy (anonymous):

but for the first question is that the answer ending at k1,k2,k3